In an aggregate expenditure model with no government or foreign sectors, represented by C = a + bY and I (an autonomous amount), saving equals investment.

Try to explain - in detail, if possible - if this is true, false, or maybe both (uncertain).
Thanks,
H.M
1
Expert's answer
2015-03-02T09:22:04-0500
In an aggregate expenditure model with no government or foreign sectors, an increase in the marginal propensity to save (s) causes the multiplier (m) to decrease, because the formula for calculating multiplier is m = 1/(1 - c) = 1/s, where s - marginal propensity to save.
That's why, if s increases, m = 1/s decreases.
So, the statement is False.